OpenStudy (anonymous):

. Find the roots of the quadratic equation x2 – 8x = 9 by completing the square. Show your work.

OpenStudy (fortytherapper):

\[x^2 - 8x = 9\] Right now it's in this form: \[ax^2-bx=c\] \[ax^2-bx+c\] ^This should be the form to get it in as a first step

OpenStudy (ibrafacts):

yes ^ follow him

OpenStudy (anonymous):

this is an easy way to do this, although you could subtract \(9 \) from both sides and then factor

OpenStudy (anonymous):

how do i factor?

OpenStudy (anonymous):

\[x^2-8x-9=0\] think of two numbers whose product is \(-9\) and that add to give \(-8\)

OpenStudy (anonymous):

i was thinking -3 but it wont add up to -8

OpenStudy (anonymous):

no \(-3\) and \(3\) add up to 0

OpenStudy (anonymous):

only other way i can think of is \(-9\) and \(1\)

OpenStudy (anonymous):

i understand now thanks :)

OpenStudy (anonymous):

yw